OpenStudy (anonymous):

P(6 2) how do you evaluate this? help please!!

sam (.sam.):

P(6 2)?

OpenStudy (anonymous):

PERMUTATIONS Yes, P(6,2)

sam (.sam.):

\[P=\frac{n!}{(n-k)!}\]

sam (.sam.):

Using 6 and 2, \[P=\frac{6!}{(6-2)!}\]

sam (.sam.):

Are you familiar with factorials?

OpenStudy (anonymous):

yes.

sam (.sam.):

So what do you get?

sam (.sam.):

6 factorial means \[6!=6 \times 5 \times 4 \times 3 \times 2 \times 1\]

OpenStudy (anonymous):

i did the equation and i get 180, what am i doing wrong?

OpenStudy (anonymous):

12 30 360 these are the answer choices.

sam (.sam.):

I'm not sure but \[\frac{6!}{(6-2)!}=\frac{6 \times 5 \times 4 \times 3 \times 2 \times 1}{4!}=\frac{6 \times 5 \times \cancel{4 \times 3 \times 2 \times 1}}{\cancel{4 \times 3 \times 2 \times 1}}=6 \times 5=30\]

sam (.sam.):

Do you get it now?

OpenStudy (anonymous):

yes, thank you! :D